# Parafermions in Moiré Minibands

### v1

@@ Abstract
Discussion.—In this work we have demonstrated the emergence of Fibonacci parafermions in the exemplary moiré system dTBG. While a quantitatively realistic material prediction would require a consideration of finite bandwidth [71] and band mixing effects [72, 73, 74] as well as reliable estimates of system parameters (which in many cases remain controversial), our results are very encouraging for several reasons.
@@
First, that an unscreened Coulomb interaction is sufficient to obtain these states is quite remarkable given that, for the simplest non-Abelian FCIs, fine-tuned interactions are necessary even in toy models  [75, 76].
@@
Second, although there is suggestive evidence for the possibility of RR parafermion quantum Hall states in Landau levels at $\nu=12/5$ and $13/5$ [58, 59], competing orders make the case less compelling than that of MR states states at $\nu=5/2$. This situation may be opposite in the moiré context: at half filling of moiré bands charge density waves [15] and the anomalous Hall crystal [77] are additional fierce competitors with no direct LL analogues, while presumably no similarly competitive non-LL states exist at $\nu=3/5$ due to a lack of natural ordering patterns.
Third, mimicking Landau levels is not the optimal strategy for finding RR parafermion states. Fig. 2 instead suggests that the most optimal bands have an average quantum geometry intermediate between the lowest and first excited Landau levels. Effects of a fluctuating geometry are unavoidable and generally favour competing states [70, 69], yet our findings suggest that the tunability of moiré materials and hence the average geometric properties are more important for the realization of RR parafermion FCIs.
@@
Fourth, although we find the RR parafermion states to be energetically favoured in a moderate parameter regime (cf. Fig. 2), it is striking that, even in cases when it is not apparent from the small system sizes avaliable whether the RR states will be competitive for a macroscopic number of particles, the entanglement spectrum shows telltale signatures of the RR states (cf. Fig. 3) suggesting that these states may prevail in a larger parameter regime.
@@
Finally, we note that Fibonacci parafermions are arguably more desirable and elusive than Majorana fermions for fundamental reasons [78]: unlike the Majoranas, they are not possible to realize in effectively non-interacting Hermitian models [52] and they potentially serve as building blocks for universal quantum computing [78]. These aspects have motivated the study of highly complex and technically extremely challenging setups such as FQH–superconductor heterostructures [51]. In contrast, moiré heterostructures offer a simple, versatile and tunable platform which does in principle not even require an external magnetic field. We hope that these observations will serve as inspiration for both the theoretical and experimental pursuit of parafermions emerging in moiré minibands.
@@ Acknowledgements.
Acknowledgements.— We acknowledge useful discussions and related collaborations with Ahmed Abouelkomsan, Liang Fu, Zhao Liu, Aidan Reddy and Donna Sheng. This work was supported by the Swedish Research Council (VR, grant 2018-00313), the Wallenberg Academy Fellows program of the Knut and Alice Wallenberg Foundation (2018.0460) and the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine.

### v2

@@ Discussion
In this work we have demonstrated the emergence of Fibonacci ($\mathbb{Z}_{3}$) parafermions in an exemplary moiré system based on dTBG. While a quantitatively realistic material prediction would require a consideration of finite bandwidth [75] and band mixing effects [76, 77, 78] as well as reliable estimates of system parameters (which in many cases remain controversial), our results are very encouraging for several reasons.
@@
First, that an unscreened Coulomb interaction is sufficient to obtain parafermions is quite remarkable given that, for the simplest non-Abelian FCIs, fine-tuned interactions are necessary even in toy models  [79, 80].
@@
Second, although there is suggestive evidence for the possibility of RR parafermion quantum Hall states in Landau levels at $\nu=12/5$ and $13/5$ [58, 59], competing orders make the case less compelling than that of MR states at $\nu=5/2$. This situation may be opposite in the moiré context: at half filling of moiré bands charge density waves [15] and the anomalous Hall crystal [81] are additional fierce competitors with no direct LL analogues, while presumably no similarly competitive non-LL states exist at $\nu=3/5$ due to a lack of natural ordering patterns.
Third, bands with an average quantum geometry equal to that of excited Landau levels are not necessarily optimal for realizing RR parafermions. Fig. 2 instead suggests that the most optimal bands have an average quantum geometry intermediate between the lowest and first excited Landau levels. It is quite likely that the optimal range of $\chi$ depends on the model/material and is related to the band flatness and quantum metric fluctuations. In this sense, the tunability of moiré bands offers enormous potential to achieve bands with the proper quantum geometric properties.
@@
Fourth, although we find the $\mathbb{Z}_{3}$ RR parafermion states to be energetically favoured in a moderate parameter range (cf. Fig. 2(a)), it is striking that, even in cases when it is not apparent from the small system sizes avaliable whether the RR states will be competitive for a macroscopic number of particles, the entanglement spectrum shows telltale signatures of the RR states (cf. Fig. 3) suggesting that these states may prevail in a larger parameter range.
@@
Finally, we emphasize that Fibonacci parafermions are arguably more desirable and elusive than Majorana fermions for fundamental reasons [82]: unlike the Majoranas, they are not possible to realize in effectively non-interacting Hermitian models [53] and they potentially serve as building blocks for universal quantum computing [82]. These aspects have motivated the study of highly complex and technically extremely challenging setups such as FQH–superconductor heterostructures [51]. In contrast, moiré heterostructures offer a simple, versatile and tunable platform which does in principle not even require an external magnetic field.
In addition, we note that although the expected Hall conductivity for the RR phase is the same as for the Abelian hierarchy state at equal filling factor, many techniques proposed for quantum Hall systems could be imported to moiré systems in order to experimentally discern these two phases. In particular, the braiding statistics of parafermions can be probed via two-point contact interferometry [83, 84, 85]. Moreover, frequency-dependent shot noise measurements in Mach-Zehnder interferometers could provide access not only to the fractional charge (which is equal in the Abelian and $\mathbb{Z}_{3}$ RR parafermion phases) but also to the non-Abelian nature of parafermions [86]. Other proposals exploit non-linear $I-V$ characteristics in tunneling experiments [87] or Coulomb blockade in quantum dots [88, 89].
We hope that these observations will serve as inspiration for both the theoretical and experimental pursuit of parafermions emerging in moiré minibands.
Methods
Quantum metric of the single-particle band
The quantum geometric properties of the target single-particle band are determined by $\phi(\mathbf{k})$, i.e. the corresponding eigenstate of the momentum-space Hamiltonian, $H_{0}(\mathbf{k})$. Here, $H_{0}(\mathbf{k})$ is a matrix in the space of layers, sublattices, and the subbands that arise from zone-folding $\mathbf{k}$-points outside of the moiré Brillouin zone up to a certain cutoff $N_{\text{cutoff}}$. Each subband element is determined by moiré reciprocal lattice vectors $\mathbf{G}$ and $\mathbf{G}^{\prime}$ and corresponds to $H_{0}(\mathbf{k})|_{\mathbf{GG}^{\prime}}=\int\mathrm{d}^{2}\mathbf{r}\ \mathrm{e}^{-i(\mathbf{k}+\mathbf{G})\cdot\mathbf{r}}H_{0}(\mathbf{r})\mathrm{e}^{i(\mathbf{k}+\mathbf{G}^{\prime})\cdot\mathbf{r}}$. The quantum (Fubini–Study) metric $g_{ij}(\mathbf{k})$ describes the distance (or form factors) between states with a small momentum difference. Concretely, it is the real part of the quantum geometric tensor $\mathcal{Q}_{ij}(\mathbf{k})$, i.e. $g_{ij}(\mathbf{k})=\mathrm{Re}[\mathcal{Q}_{ij}(\mathbf{k})]$ where $\mathcal{Q}_{ij}(\mathbf{k})=\partial_{i}\phi^{\dagger}(\mathbf{k})\partial_{j}\phi(\mathbf{k})-[\partial_{i}\phi^{\dagger}(\mathbf{k})\phi(\mathbf{k})][\phi^{\dagger}(\mathbf{k})\partial_{j}\phi(\mathbf{k})]$ ($\partial_{i}\equiv\partial_{k_{i}}$ with $i=x,y$).
Exact diagonalization
In order to tackle the many-body problem of interacting electrons, we first diagonalize $H_{0}(\bf r)$ in the basis of Bloch states and then project the electron-electron interactions onto the considered nearly-flat band, obtaining $H_{\text{int}}=\frac{1}{2}\sum_{\bf k_{1}k_{2}k_{3}k_{4}}V_{\bf k_{1}k_{2}k_{3}k_{4}}c_{\bf k_{1}}^{\dagger}c_{\bf k_{2}}^{\dagger}c_{\bf k_{3}}c_{\bf k_{4}}.$ Here, the matrix element $V_{\bf k_{1}k_{2}k_{3}k_{4}}$ ensures momentum conservation and contains single-particle form factors [90, 8] as well as the 2D bare Coulomb potential $V({\bf q})=\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|{\bf q}|}$ with the average dielectric constant of the system $\epsilon=5$ [91]. Next, we obtain the low-lying energy spectrum of many-body states by exact diagonalization of the spin- and valley-polarized interaction Hamiltonian $H_{\text{int}}$. In particular, we consider a system of $N_{\text{e}}$ electrons in $N_{\text{s}}$ moiré sites with the filling $\nu=N_{\text{e}}/N_{\text{s}}=3/5,2/5$. We ensure that the calculations are converged with respect to the number of moiré subbands, $(2N_{\text{cutoff}}+1)^{2}$, and note that $N_{\text{cutoff}}=4$ is sufficient. In addition, we note that calculations in finite-size systems yield results that deviate with respect to infinite systems due to finite size effects. Here, we have employed state of the art methods and considered systems as large as possible to minimize such uncertainties.
Structure factor and pair correlation
The structure factor helps measure if the ground states exhibit a preferred charge order and is defined as $S(\mathbf{q})=\frac{1}{N_{\text{e}}^{2}}\langle\rho(\mathbf{q})\rho(-\mathbf{q})\rangle-\delta_{\mathbf{q},0},$ with $\rho(\mathbf{q})$ being the density operator projected onto the targeted nearly flat band. On the other hand, the real space pair-correlation function reads $C(\mathbf{r})=\frac{1}{N^{2}_{\text{e}}}\braket{n(\mathbf{r})n(\mathbf{0})}$, where $n(\mathbf{r})$ is the real-space density operator. CDW order that breaks the translation symmetry of the underlying moiré lattice is characterized by structure factor peaks at symmetry points of the Brillouin zone, whereas liquids are distinguished by the absence of such peaks. Symmetry-breaking CDW peaks in $S(\mathbf{q})$ would correspond to a periodic modulation of the pair-correlation function with a periodicity different than that of the underlying moiré lattice.
Data availability
The data generated in this study and presented in the article are provided in the Source Data file.
Code availability
The codes used to generate and analyse the data are available from the corresponding author upon request.
@@ Acknowledgements.
Acknowledgements
We acknowledge useful discussions and related collaborations with Ahmed Abouelkomsan, Liang Fu, Zhao Liu, Aidan Reddy and Donna Sheng. This work was supported by the Swedish Research Council (VR, grant 2018-00313), the Wallenberg Academy Fellows program of the Knut and Alice Wallenberg Foundation (2018.0460) and the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Author contributions
E.B. conceived the research. H.L. and R.P.-C. performed the numerical calculations. All authors analyzed and discussed the results and wrote the manuscript.
Competing interests
The authors declare that they have no competing interests.